OpenStudy (anonymous):

Solve the equation in the interval [0, 2pi]: 2sin^2x-5sinx+2=0

OpenStudy (anonymous):

\[\Large 2 \sin ^2(x)-5 \sin (x)+2=(\sin (x)-2) (2 \sin (x)-1)=0\\ \Large \sin(x)=\frac 1 2\\ \sin(x)=2\quad \text {( Impossible)} \]

OpenStudy (anonymous):

\[ \Large \sin(x)=\frac 1 2\implies x=\frac \pi 6 + 2 k \pi \text{ or } x=\frac {5\pi}6 + 2 k \pi \] Choose the solutions in the given interval

OpenStudy (whpalmer4):

For the actual solving, it may be easier to do it if you make a substitution first: let \(u = \sin x\) \[2\sin^2(x) - 5\sin(x) + 2 =0\]\[ 2u^2-5u+2=0\](solve by your favorite method) \[u = 2, \,u=\frac{1}{2}\]Undo the substitution: \[\sin(x) = 2\](impossible, as previously mentioned, so disregard) \[\sin(x) = \frac{1}{2}\]\[x = \frac{\pi}{6},\,\frac{5\pi}{6}\]are the solutions in the specified interval.